# SCMM: Calibrating Cross-modal Representations for Text-Based Person Search

### v5

@@ III-A Sew Calibration Loss with Constraints
As shown in Fig. 3(a), in single modality settings, a triplet distance constraint is widely utilized when embeddings are from a single modality distribution. In intra-class samples, the embeddings are pulled together, while inter-class samples are pushed away. In the cross-modal setting, we expect to impose constraints from both sides of the two embedding distributions. For example, in one-direction retrieval, we first need to align the features of a ”perfect pair,” i.e., $(I_{i},T_{i})$. The shortest distance from the image embedding ${v_{i}}$ to the text embedding should be ${t_{i}}$. This is because it is a perfect matching pair in person re-identification, and no other text feature will have a shorter distance (Eq. 1). We set the perfect pair as an image-text anchor $(A_{img},A_{txt})$. Next, we impose another constraint between the image anchor $A_{img}$ and its corresponding positive text samples $P_{txt}(\mathds{1}(y_{i}=y_{j},i\neq j))$ and negative text samples $N_{txt}$ (Eq. 2). For the other direction, we put symmetric constraints on $A_{txt}$ and $P_{img}$, $N_{img}$. Fig. 3(b) shows the proposed constraints for cross modality. Forces from both sides act like a seam to pull the two distributions together.
@@
III-B Masking Caption Modeling Loss
TBPS is a fine-grained cross-modal task, which means only caption-level discrimination is not enough. If the textual captions of two persons differ in a few words, a TBPS method can not retrieve a specific person without word-level discrimination. Although there are many works to establish word-level discrimination capacities, such methods are complex and limited, hindering performance boost. To solve this issue, we propose masking caption modeling to establish detailed image-text relationships. Furthermore, by utilizing MCM, our framework can perform more generic cross-modal learning.
@@ III-B Masking Caption Modeling Loss
Input: Image $I$ and text $T$; A batch of $n$ paired $\mathbb{G}=\{(I^{1},T^{1}),(I^{2},T^{2}),\cdots,(I^{n},T^{n})\}$
@@
1 foreach $T^{i}$ in set $\mathbb{G}$ do
2      $T^{i}\leftarrow\text{mask}(T^{i})$;
3 end foreach
4for $i\leftarrow 1$ to $\mathbb{G}$ do
5       $(I^{i}_{cls},I^{i}_{1},\cdots,I^{i}_{c})\leftarrow ViT(T^{i})$;
6       $(T^{i}_{cls},T^{i}_{1},\cdots,T^{i}_{c})\leftarrow Bert(T^{i})$;
7       if Training Stage then
8             $(I^{i}_{cls},T^{i}_{cls})\leftarrow BatchNorm(I^{i}_{cls},T^{i}_{cls})$;
9             $I^{i}_{context},T^{i}_{context}\leftarrow(I^{i}_{1},\cdots,I^{i}_{c}),(T^{i}_{1},\cdots,T^{i}_{c})$;
10             $T^{i}_{context}\leftarrow Attn(T^{i}_{context})$;
11             $T^{i}_{cross}\leftarrow CrossAttn(T^{i}_{context},I^{i}_{context})$;
12             $\mathcal{L}_{sew}^{i}\leftarrow SewCalibration(I^{i}_{cls},T^{i}_{cls})$;
13             $\mathcal{L}_{mcm}^{i}\leftarrow MCM(T^{i}_{cross},T^{i})$;
14
15      else
16             $\{(I^{1}_{cls},T^{1}_{cls}),(I^{2}_{cls},T^{2}_{cls}),\cdots,(I^{i}_{cls},T^{i}_{cls})\}\leftarrow(I_{cls},T_{cls})$;
17
18       end if
19
20 end for
21if Training Stage then
22       return $(\mathcal{L}_{sew},\mathcal{L}_{mcm})$;
23
24else
25       return $(I_{cls},T_{cls})$;
26
27 end if
28
Algorithm 1 Traning Procedure of SCMM
@@ 13             $\mathcal{L}_{mcm}^{i}\leftarrow MCM(T^{i}_{cross},T^{i})$;
III-C Total Loss
The total loss $\mathcal{L}$ we optimized in each iteration is as follows:
$\mathcal{L}_{sew}=\mathcal{L}_{match}^{I2T}+\mathcal{L}_{match}^{T2I}+\mathcal{L}_{id}^{I2T}+\mathcal{L}_{id}^{T2I},$
@@ III-C Total Loss
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters to balance the different loss terms during training. The pseudocode of our framework pipeline is shown in Algorithm. 1.
IV Experiments
IV-A Datasets and Evaluation Metric
@@
We evaluate SCMM on three benchmarks: CUHK-PEDES [10], ICFG-PEDES [19], and RSTPReid [26]. CUHK-PEDES is the first large-scale benchmark for text-based person search tasks. This dataset contains 40,206 images of 13,003 person IDs collected from five person re-identification datasets. Each image has two different textual captions with an extensive vocabulary, and the average sentence length is 23.5. The testing set comprises 3,074 images and 6,148 descriptions of 1,000 persons. ICFG-PEDES contains 54,522 images of 4,102 persons. For each image, the corresponding description has an average length of 37 words. The testing set consists of 19,848 image-text pairs of 1,000 persons. RSTPReID contains 20,505 images of 4,101 persons, with each pedestrian having five images. It is divided into a training set with 3,701 persons, a validation set with 200 persons, and a testing set with 200 persons. For our evaluation metric, we report the Rank $k$ ($k$=1, 5, 10) text-to-image accuracy, which is commonly used in previous works to evaluate text-based person search. Given a textual description as the query, if the top-$k$ retrieved images contain any person corresponding to the query, we consider it a successful person search.
IV-B Implementation Details
In our visual-textual dual-encoder, we extract visual representations using the ViT-Base pre-trained on ImageNet [42]. The images are resized to 224 $\times$ 224 pixels. For textual representations, we use the BERT-Base-Uncased model pre-trained on the Toronto Book Corpus and Wikipedia. The representation dimension is set to 768, and the feature sequence lengths are set to 197 and 100, respectively.
@@ IV-B Implementation Details
During the training phase, we use a batch size of 64 and train for 60 epochs. We use Adam optimizer with an initial learning rate of 0.001. To augment our data, we apply a random horizontal flipping operation, and we use a mask ratio of 0.1 for randomly masking text tokens. The minimum and maximum textual information length boundaries $\mathcal{T}_{min}$ and $\mathcal{T}_{max}$ are set to 20-60, 25-65, and 22-60 according to the caption length distributions of CUHK-PEDES, ICFG-PEDES, and RSTPReID, respectively. The bounds of the upper and lower margin $\mathcal{M}$ are set to 0.4 and 0.6, and the scale parameter $\alpha$ is set to 32. For each loss in the total loss function, the balance factors $\lambda_{1}$ and $\lambda_{2}$ are set equal to 1. During the testing phase, we apply a re-ranking post-processing approach to improve search performance following NAFS [28]. We conduct the experiments on four NVIDIA 2080Ti GPUs using PyTorch.
@@
TABLE I: Comparison with SOTA methods on the CUHK-PEDES dataset. Rank1, Rank5, and Rank10 accuracies (%) are reported.
@@
  • Bold number represents the best score. R and C denote the re-ranking post-processing operations and image encoder pre-trained on CLIP model.
TABLE II: Quantitative results on the ICFG-PEDES dataset.
@@
IV-C Comparison with State-of-the-art Methods
IV-C1 Results on CUHK-PEDES.
Table I compares our framework and previous methods on CUHK-PEDES. It can be observed that our proposed SCMM can outperform all previous methods by a large margin. Compared to the state-of-the-art work CFine [24], SCMM achieves 67.71%(+2.64%), 84.57%(+1.56%) and 89.44%(+0.44%) on Rank1, Rank5 and Rank10 without re-ranking. The state-of-the-art results on the CUHK-PEDES show the effectiveness of SCMM. With the help of re-ranking, SCMM shows an incremental boost to get a 71.09% Rank1 score. Furthermore, with a better image encoder pre-training on CLIP, SCMM achieves 73.81%, 88.89% , and 92.77% on three metrics, respectively. These consistent improvements show the scalability of SCMM across better pre-trained models and extra post-processing operations.
IV-C2 Results on ICFG-PEDES and RSTPReid.
We also utilized other benchmarks to validate SCMM’s performance and generalization. The ICFG-PEDES and RSTPReid datasets are more challenging compared to CUHK-PEDES, and our method significantly outperformed all state-of-the-art methods on these two datasets by a large margin, as reported in Tables II and III. Compared with the state-of-the-art results [24] on ICFG-PEDES, SCMM achieved significant improvements, with scores of 60.20%(+4.51%), 75.97%(+3.25%), and 81.78%(+2.32%) on Rank1, Rank5, and Rank10, respectively. On the RSTPReid dataset, we achieved scores of 50.75% (+4.90%), 74.20%(+3.90%), and 81.70%(+3.30%) on the three metrics. With re-ranking post-processing, we were able to achieve scores of 62.17% and 55.35% on Rank1 for the two benchmarks, respectively.
We note that re-ranking also brings a significant boost, as our model learns clear and compact cross-modal key information patterns, and re-ranking can retrieve the correct feature neighbors more effectively. Moreover, by utilizing the CLIP pre-trained image model as an encoder, we achieved even better results, with scores of 64.25%, 86.95%, and 90.70% on ICFG-PEDES, and scores of 57.35%, 77.50%, and 85.50% on RSTPReid. These two challenging benchmark results demonstrate the robustness and scalability of SCMM.
TABLE III: Quantitative results on the RSTPReid dataset.
@@ IV-C Comparison with State-of-the-art Methods
IV-D Ablation Studies
IV-D1 Analysis of Model Components
To fully validate the performance of the different components in SCMM, we demonstrate the contributions of each part on CUHK-PEDES, as shown in Table IV. Model 1 and Model 2 show the results using ResNet [51] and ViT [39] as the image encoder, respectively, without the Sew Calibration and MCM losses. We utilize Model 2 as our baseline, and CMPM and CMPC [14] are used as loss functions in the baseline experiment. First, we compare ResNet-50 and ViT-Base as the image encoder and observe a performance improvement of 1.90%, 1.03%, and 0.69% on Rank1, Rank5, and Rank10, respectively. Based on this, we adopt ViT as the image encoder in the following experiments. We then validate the Sew calibration and MCM losses compared to the baseline. From Models 2-4, we observe that Sew calibration loss brings marginal improvement, and a common distribution can provide a better basis for optimizing the embedding space for fine-grained cross-modal recognition. Moreover, Model 3 represents the Sew calibration loss with the fixed margin 0.5, while Model 4 represents the Sew calibration loss with adaptive margins. We observe that Model 4 achieves better results than a fixed manual margin, no matter how we tune the margin value, which demonstrates the effectiveness of adaptive constraints. We also find that MCM gives another substantial improvement to the baseline from Model 2 to Model 5, indicating that the text-image detail mining capability is critical. Notably, when the two components are used jointly in Model 6, SCMM continues to improve and outperform the baseline by 4.57%, 2.99%, and 2.37% on the three metrics, respectively. This demonstrates that obtaining well-distributed image-text features in the common embedding space is essential for reducing the cross-modal gap. The consistent improvement in each component of the method demonstrates our effectiveness.
TABLE IV: Performance comparisons of different components.
@@ IV-D1 Analysis of Model Components
  • Res and ViT denote ResNet and Vision Transformer as image encoders,
  • respectively. Sew-F and Sew-A refer to the fixed and adaptive margin
  • versions of the Sew calibration loss, respectively. Model 2 is our baseline.
IV-D2 Impact of Sew Calibration Loss for Reducing Cross-modal Gap
@@
TABLE V: MCM components on the CUHK-PEDES dataset.

### v7

@@ 3.1 Sew Calibration Loss with Constraints
In the TBPS task, closely aligning cross-modal representations is crucial for effectively finding specific persons with textual captions. To address the heterogeneity between modalities, we propose a sew calibration loss that pushes each modality to a common space. As shown in Fig. 3(a), in single modality settings, a triplet distance constraint is widely utilized when embeddings are from a single modality distribution. In intra-class samples, the embeddings are pulled together, while inter-class samples are pushed away. In the cross-modal setting, we expect to impose constraints from both sides of the two embedding distributions. For example, in one-direction retrieval, we first need to align the features of a ”perfect pair,” i.e., $(I_{i},T_{i})$. The shortest distance from the image embedding ${v_{i}}$ to the text embedding should be ${t_{i}}$. Consequently, it is a perfect matching pair in person re-identification, and no other text feature will have a shorter distance (Eq. 1). We set the perfect pair as an image-text anchor $(A_{img},A_{txt})$. Next, we impose another constraint between the image anchor $A_{img}$ and its corresponding positive text samples $P_{txt}(\mathds{1}(y_{i}=y_{j},i\neq j))$ and negative text samples $N_{txt}$ (Eq. 2). For the other direction, we put symmetric constraints on $A_{txt}$ and $P_{img}$, $N_{img}$. Fig. 3(b) shows the proposed constraints for cross modality. Forces from both sides act like a seam to pull the two distributions together.
(a) Single-modal constraints
(b) Cross-modal constraints
@@
3.2 Masked Caption Modeling Loss
TBPS is a fine-grained cross-modal task, which means only caption-level discrimination is not enough. If the textual captions of two persons differ in a few words, a TBPS method can not retrieve a specific person without word-level discrimination. Although there are many works to establish word-level discrimination capacities, such methods are complex and limited, hindering performance boost. To solve this issue, we propose masked caption modeling to establish detailed image-text relationships. Furthermore, by utilizing MCM, our framework can perform more generic cross-modal learning.
@@ 3.2 Masked Caption Modeling Loss
Input:
Image $I$ and text $T$; A batch of $n$ paired $\mathbb{G}=\{(I^{1},T^{1}),(I^{2},T^{2}),\cdots,(I^{n},T^{n})\}$
@@
foreach $T^{i}$ in set $\mathbb{G}$ do
      $T^{i}\leftarrow\text{mask}(T^{i})$;
end foreach
for $i\leftarrow 1$ to $\mathbb{G}$ do
       $(I^{i}_{cls},I^{i}_{1},\cdots,I^{i}_{c})\leftarrow ViT(I^{i})$;
       $(T^{i}_{cls},T^{i}_{1},\cdots,T^{i}_{c})\leftarrow Bert(T^{i})$;
       if Training Stage then
             $(I^{i}_{cls},T^{i}_{cls})\leftarrow BatchNorm(I^{i}_{cls},T^{i}_{cls})$;
             $I^{i}_{context},T^{i}_{context}\leftarrow(I^{i}_{1},\cdots,I^{i}_{c}),(T^{i}_{1},\cdots,T^{i}_{c})$;
             $T^{i}_{context}\leftarrow Attn(T^{i}_{context})$;
             $T^{i}_{cross}\leftarrow CrossAttn(T^{i}_{context},I^{i}_{context})$;
             $\mathcal{L}_{sew}^{i}\leftarrow SewCalibration(I^{i}_{cls},T^{i}_{cls})$;
             $\mathcal{L}_{mcm}^{i}\leftarrow MCM(T^{i}_{cross},T^{i})$;
      else
             $\{(I^{1}_{cls},T^{1}_{cls}),(I^{2}_{cls},T^{2}_{cls}),\cdots,(I^{i}_{cls},T^{i}_{cls})\}\leftarrow(I_{cls},T_{cls})$;
       end if
end for
if Training Stage then
       return $(\mathcal{L}_{sew},\mathcal{L}_{mcm})$;
else
       return $(I_{cls},T_{cls})$;
end if
Algorithm 1 Training Procedure of SCMM
@@
3.3 Total Loss
The total loss function of SCMM integrates our two complementary fusion mechanisms to achieve comprehensive cross-modal alignment and correspondence learning. Our unified optimization objective combines the sew calibration loss and the masked caption modeling loss to jointly address both global alignment and fine-grained correspondence challenges in cross-modal information fusion. The sew calibration loss $\mathcal{L}_{sew}$ encompasses bidirectional matching and classification components that ensure symmetric cross-modal alignment:
$\mathcal{L}_{sew}=\mathcal{L}_{match}^{I2T}+\mathcal{L}_{match}^{T2I}+\mathcal{L}_{id}^{I2T}+\mathcal{L}_{id}^{T2I},$ (12)
where the image-to-text and text-to-image matching losses establish cross-modal correspondence constraints, while the identification losses enforce discriminative learning for person-specific features. Importantly, the bidirectional formulation ensures that both modalities are effectively aligned in the shared embedding space through symmetric constraints. Subsequently, the complete loss function combines both components through weighted summation:
@@ 3.3 Total Loss
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters balancing global alignment and fine-grained correspondence learning. Through this joint optimization, the framework simultaneously learns compact cross-modal representations while capturing detailed visual-textual relationships for discriminating similar individuals. The complementary loss components create synergistic effects ensuring comprehensive information fusion, addressing both inter-modal gap reduction and fine-grained correspondence modeling. Regarding computational complexity, SCMM maintains efficient processing through its streamlined dual-encoder architecture. Training complexity is $O(n\cdot d^{2})$ where $n$ is batch size and $d$ is feature dimension, dominated by cross-attention operations in the MCM decoder. During inference, only CLS token embeddings are utilized, achieving $O(d)$ complexity per query-gallery comparison. In contrast, the method contrasts favorably with attention-explicit methods like ACSA [26] requiring $O(n^{2}\cdot d)$ complexity due to full cross-modal attention during inference. Our training-only decoder design ensures high-speed retrieval while maintaining sophisticated cross-modal learning capabilities.
4 Experiments
4.1 Datasets and Evaluation Metric
@@
We evaluate SCMM on three benchmarks: CUHK-PEDES [8], ICFG-PEDES [20], and RSTPReid [13]. CUHK-PEDES is the first large-scale benchmark for text-based person search tasks. As the foundational dataset, it contains 40,206 images of 13,003 person IDs collected from five person re-identification datasets. Each image has two different textual captions with an extensive vocabulary, and the average sentence length is 23.5. The testing set comprises 3,074 images and 6,148 descriptions of 1,000 persons. ICFG-PEDES contains 54,522 images of 4,102 persons. For each image, the corresponding description has an average length of 37 words. The testing set consists of 19,848 image-text pairs of 1,000 persons. RSTPReID contains 20,505 images of 4,101 persons, with each pedestrian having five images. It is divided into a training set with 3,701 persons, a validation set with 200 persons, and a testing set with 200 persons. For our evaluation metric, we report the Rank $k$ ($k$=1, 5, 10) text-to-image accuracy, which is commonly used in previous works to evaluate text-based person search. Given a textual description as the query, if the top-$k$ retrieved images contain any person corresponding to the query, we consider it a successful person search.
4.2 Implementation Details
@@
In our visual-textual dual-encoder, we extract visual representations using the ViT-Base pre-trained on ImageNet [38]. The images are resized to 224 $\times$ 224 pixels. For textual representations, we use the BERT-Base-Uncased model pre-trained on the Toronto Book Corpus and Wikipedia. The representation dimension is set to 768, and the feature sequence lengths are set to 197 and 100, respectively. During the training phase, we use a batch size of 64 and train for 60 epochs. We use Adam optimizer with an initial learning rate of 0.001. To augment our data, we apply a random horizontal flipping operation, and we use a mask ratio of 0.1 for randomly masking text tokens. The minimum and maximum textual information length boundaries $\mathcal{T}_{min}$ and $\mathcal{T}_{max}$ are set to 20-60, 25-65, and 22-60 according to the caption length distributions of CUHK-PEDES, ICFG-PEDES, and RSTPReID, respectively. The bounds of the upper and lower margin $\mathcal{M}$ are set to 0.4 and 0.6, and the scale parameter $\alpha$ is set to 32. For each loss in the total loss function, the balance factors $\lambda_{1}$ and $\lambda_{2}$ are set equal to 1. During the testing phase, we apply a re-ranking post-processing approach to improve search performance following NAFS [14]. We conduct the experiments on four NVIDIA 2080Ti GPUs using PyTorch.
4.3 Comparison with State-of-the-art Methods
@@
Table 1: Comparison with SOTA methods on the CUHK-PEDES dataset. Rank1, Rank5, and Rank10 accuracies (%) are reported. Bold number represents the best score. R and C denote the re-ranking post-processing operations and image encoder pre-trained on CLIP model.
@@
4.3.1 Results on CUHK-PEDES.
Table 1 compares our framework and previous methods on CUHK-PEDES. It can be observed that our proposed SCMM can outperform all previous methods by a large margin. Compared to the work CFine [25], SCMM achieves 67.71% (+2.64%), 84.57% (+1.56%) and 89.44% (+0.44%) on Rank1, Rank5 and Rank10 without re-ranking. The state-of-the-art results on the CUHK-PEDES show the effectiveness of SCMM. With the help of re-ranking, SCMM shows an incremental boost to get a 71.09% Rank1 score. Furthermore, with a better image encoder pre-training on CLIP, SCMM achieves 73.81%, 88.89% , and 92.77% on three metrics, respectively. Overall, the consistent improvements show the scalability of SCMM across better pre-trained models and extra post-processing operations.
Table 2: Quantitative results on the ICFG-PEDES dataset.
@@
Table 3: Quantitative results on the RSTPReid dataset.
@@
4.3.2 Results on ICFG-PEDES and RSTPReid.
We also utilized other benchmarks to validate SCMM’s performance and generalization. The ICFG-PEDES and RSTPReid datasets are more challenging compared to CUHK-PEDES, and our method significantly outperformed all state-of-the-art methods on these two datasets by a large margin, as reported in Tables 2 and 3. Compared with the state-of-the-art results [25] on ICFG-PEDES, SCMM achieved significant improvements, with scores of 60.20% (+4.51%), 75.97% (+3.25%), and 81.78% (+2.32%) on Rank1, Rank5, and Rank10, respectively. On the RSTPReid dataset, we achieved scores of 50.75% (+4.90%), 74.20% (+3.90%), and 81.70% (+3.30%) on the three metrics. With re-ranking post-processing, we were able to achieve scores of 62.17% and 55.35% on Rank1 for the two benchmarks, respectively.
We note that re-ranking also brings a significant boost, as our model learns clear and compact cross-modal key information patterns, and re-ranking can retrieve the correct feature neighbors more effectively. Moreover, by utilizing the CLIP pre-trained image model as an encoder, we achieved even better results, with scores of 64.25%, 86.95%, and 90.70% on ICFG-PEDES, and scores of 57.35%, 77.50%, and 85.50% on RSTPReid. Consequently, the two challenging benchmark results demonstrate the robustness and scalability of SCMM.
4.4 Ablation Studies
Table 4: Performance comparisons of different components. Res and ViT denote ResNet and Vision Transformer as image encoders, respectively. Sew-F and Sew-A refer to the fixed and adaptive margin versions of the sew calibration loss, respectively. Model 2 is our baseline.
@@ 4.4 Ablation Studies
4.4.1 Analysis of Model Components
To fully validate the performance of the different components in SCMM, we demonstrate the contributions of each part on CUHK-PEDES, as shown in Table 4. Model 1 and Model 2 show the results using ResNet [42] and ViT [37] as the image encoder, respectively, without the Sew Calibration and MCM losses. We utilize Model 2 as our baseline, and CMPM and CMPC [16] are used as loss functions in the baseline experiment. First, we compare ResNet-50 and ViT-Base as the image encoder and observe a performance improvement of 1.90%, 1.03%, and 0.69% on Rank1, Rank5, and Rank10, respectively. Based on this, we adopt ViT as the image encoder in the following experiments. We then validate the sew calibration and MCM losses compared to the baseline. From Models 2-4, we observe that sew calibration loss brings marginal improvement, and a common distribution can provide a better basis for optimizing the embedding space for fine-grained cross-modal recognition. Moreover, Model 3 represents the sew calibration loss with the fixed margin 0.5, while Model 4 represents the sew calibration loss with adaptive margins. We observe that Model 4 achieves better results than a fixed manual margin, no matter how we tune the margin value, which demonstrates the effectiveness of adaptive constraints. We also find that MCM gives another substantial improvement to the baseline from Model 2 to Model 5, indicating that the text-image detail mining capability is critical. Notably, when the two components are used jointly in Model 6, SCMM continues to improve and outperform the baseline by 4.57%, 2.99%, and 2.37% on the three metrics, respectively. Importantly, the results demonstrate that obtaining well-distributed image-text features in the common embedding space is essential for reducing the cross-modal gap. The consistent improvement in each component of the method demonstrates our effectiveness.
4.4.2 Impact of Sew Calibration Loss for Reducing Cross-modal Gap
@@
4.4.3 Impact of Masked Caption Modeling
The masked caption modeling operation in fine-grained cross-modal interaction is critical in our framework. It comprises tokens masking, attention module, and masked tokens prediction. As shown in Table 5, we explore the effectiveness of MCM components on the CUHK-PEDES. First, only masking on the input text tokens without a reconstruction task behaves like a random erase text augmentation. Remarkably, the augmentation can already bring +1.36% marginal improvement and reach 63.65% on Rank1. It shows that the details provided by word tokens are helpful for fine-grained recognition.
Table 5: MCM components on the CUHK-PEDES dataset.